# Impact of 3D macro physics and nuclear physics on the $p$ nuclei in $\mathrm{O}$-$\mathrm{C}$ shell mergers

### v1

@@ Appendix C Results of varying the input nuclear reactions
${}^{74}\mathrm{Se}$ ${}^{75}\mathrm{Se}(\gamma,n)$ $-0.83$ $-0.22$ ${}^{152}\mathrm{Gd}$ ${}^{150}\mathrm{Gd}(\gamma,\alpha)$ $-0.18$ $-0.00$
${}^{78}\mathrm{Kr}$ ${}^{79}\mathrm{Kr}(\gamma,n)$ $-0.79$ $-0.24$ ${}^{151}\mathrm{Gd}(\gamma,n)$ $-0.20$ $-0.00$
${}^{84}\mathrm{Sr}$ ${}^{85}\mathrm{Sr}(\gamma,n)$ $-0.80$ $-0.34$ ${}^{152}\mathrm{Dy}(\gamma,\alpha)$ $-0.31$ $-0.01$
${}^{92}\mathrm{Mo}$ ${}^{93}\mathrm{Mo}(\gamma,n)$ $-0.91$ $-0.14$ ${}^{154}\mathrm{Dy}(\gamma,\alpha)$ $-0.17$ $-0.00$
@@
${}^{96}\mathrm{Ru}$ ${}^{97}\mathrm{Ru}(\gamma,n)$ $-0.82$ $-0.14$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.00$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $0.21$ $0.04$ ${}^{156}\mathrm{Dy}$ ${}^{157}\mathrm{Dy}(\gamma,n)$ $-0.28$ $-0.12$
${}^{98}\mathrm{Ru}$ ${}^{97}\mathrm{Ru}(\gamma,n)$ $0.50$ $0.02$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.17$ $-0.05$
@@
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.31$ $-0.01$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.05$
@@
${}^{102}\mathrm{Pd}$ ${}^{100}\mathrm{Pd}(\gamma,p)$ $-0.18$ $-0.04$ ${}^{159}\mathrm{Dy}(\gamma,n)$ $-0.21$ $-0.03$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.21$ $-0.04$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.17$ $-0.02$
${}^{103}\mathrm{Pd}(\gamma,n)$ $-0.72$ $-0.16$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $0.66$ $0.07$
${}^{106}\mathrm{Cd}$ ${}^{107}\mathrm{Cd}(\gamma,n)$ $-0.75$ $-0.16$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.02$
@@
${}^{108}\mathrm{Cd}$ ${}^{107}\mathrm{Cd}(\gamma,n)$ $0.20$ $0.05$ ${}^{162}\mathrm{Er}$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.25$ $-0.10$
${}^{109}\mathrm{Cd}(\gamma,n)$ $-0.83$ $-0.20$ ${}^{159}\mathrm{Er}(\gamma,\alpha)$ $-0.16$ $-0.05$
${}^{113}\mathrm{In}$ ${}^{114}\mathrm{In}(\gamma,n)$ $-0.32$ $-0.05$ ${}^{160}\mathrm{Er}(\gamma,n)$ $-0.23$ $-0.09$
${}^{110}\mathrm{Sn}(\gamma,p)$ $-0.16$ $-0.02$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $-0.44$ $-0.15$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.27$ $-0.04$ ${}^{161}\mathrm{Er}(\gamma,n)$ $0.16$ $0.06$
@@
${}^{112}\mathrm{Sn}$ ${}^{113}\mathrm{Sn}(\gamma,n)$ $-0.80$ $-0.18$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.06$
${}^{114}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.17$ $-0.01$ ${}^{164}\mathrm{Er}$ ${}^{164}\mathrm{Yb}(\gamma,n)$ $-0.26$ $-0.07$
${}^{113}\mathrm{Sn}(\gamma,n)$ $0.65$ $0.05$ ${}^{164}\mathrm{Yb}(\gamma,\alpha)$ $-0.56$ $-0.18$
@@
${}^{115}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.16$ $-0.01$ ${}^{168}\mathrm{Yb}$ ${}^{164}\mathrm{Yb}(\gamma,\alpha)$ $-0.19$ $-0.09$
${}^{113}\mathrm{Sn}(\gamma,n)$ $0.74$ $0.06$ ${}^{166}\mathrm{Yb}(\gamma,n)$ $-0.22$ $-0.08$
${}^{122}\mathrm{Xe}(\gamma,p)$ $0.16$ $0.01$ ${}^{166}\mathrm{Yb}(\gamma,\alpha)$ $-0.38$ $-0.13$
@@
${}^{120}\mathrm{Te}$ ${}^{121}\mathrm{Te}(\gamma,n)$ $-0.76$ $-0.19$ ${}^{172}\mathrm{Hf}(\gamma,\alpha)$ $0.43$ $0.16$
@@
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.23$ $-0.05$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.19$ $0.06$
${}^{124}\mathrm{Xe}$ ${}^{122}\mathrm{Xe}(\gamma,p)$ $-0.19$ $-0.05$ ${}^{174}\mathrm{Hf}$ ${}^{172}\mathrm{Hf}(\gamma,\alpha)$ $-0.40$ $-0.07$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.39$ $-0.15$ ${}^{174}\mathrm{W}(\gamma,\alpha)$ $-0.19$ $-0.04$
${}^{125}\mathrm{Xe}(\gamma,n)$ $-0.54$ $-0.23$ ${}^{178}\mathrm{W}(\gamma,\alpha)$ $0.51$ $0.09$
${}^{126}\mathrm{Xe}$ ${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $-0.35$ $-0.14$ ${}^{182}\mathrm{Os}(\gamma,\alpha)$ $0.18$ $0.03$
@@
${}^{127}\mathrm{Xe}(\gamma,n)$ $-0.63$ $-0.27$ ${}^{180}\mathrm{Ta}$ ${}^{179}\mathrm{Ta}(\gamma,n)$ $-0.91$ $-0.07$
${}^{130}\mathrm{Ba}$ ${}^{131}\mathrm{Ba}(\gamma,n)$ $-0.82$ $-0.13$ ${}^{179}\mathrm{Ta}(\gamma,\alpha)$ $-0.15$ $-0.01$
${}^{132}\mathrm{Ba}$ ${}^{131}\mathrm{Ba}(\gamma,n)$ $0.36$ $0.14$ ${}^{180}\mathrm{W}$ ${}^{180}\mathrm{Os}(\gamma,n)$ $-0.27$ $-0.05$
${}^{133}\mathrm{Ba}(\gamma,n)$ $-0.76$ $-0.30$ ${}^{180}\mathrm{Os}(\gamma,\alpha)$ $-0.51$ $-0.13$
${}^{138}\mathrm{La}$ ${}^{133}\mathrm{La}(\gamma,p)$ $-0.20$ $-0.07$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.23$ $0.04$
${}^{135}\mathrm{La}(\gamma,n)$ $-0.35$ $-0.12$ ${}^{184}\mathrm{Os}$ ${}^{182}\mathrm{Os}(\gamma,\alpha)$ $-0.38$ $-0.02$
${}^{136}\mathrm{La}(\gamma,n)$ $0.26$ $0.08$ ${}^{184}\mathrm{Pt}(\gamma,\alpha)$ $-0.33$ $-0.02$
${}^{137}\mathrm{La}(\gamma,n)$ $-0.34$ $-0.14$ ${}^{186}\mathrm{Pt}(\gamma,\alpha)$ $-0.15$ $-0.01$
${}^{136}\mathrm{Ce}$ ${}^{137}\mathrm{Ce}(\gamma,n)$ $-0.49$ $-0.07$ ${}^{188}\mathrm{Pt}(\gamma,\alpha)$ $0.23$ $0.01$
${}^{138}\mathrm{Nd}(\gamma,n)$ $-0.27$ $-0.03$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.19$ $0.01$
${}^{138}\mathrm{Nd}(\gamma,p)$ $0.48$ $0.06$ ${}^{190}\mathrm{Pt}$ ${}^{190}\mathrm{Hg}(\gamma,n)$ $-0.30$ $-0.02$
${}^{140}\mathrm{Nd}(\gamma,\alpha)$ $0.38$ $0.04$ ${}^{190}\mathrm{Hg}(\gamma,\alpha)$ $-0.50$ $-0.05$
@@
${}^{139}\mathrm{Ce}(\gamma,n)$ $-0.52$ $-0.05$ ${}^{196}\mathrm{Hg}$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $-0.68$ $-0.35$
${}^{138}\mathrm{Nd}(\gamma,p)$ $0.16$ $0.01$ ${}^{197}\mathrm{Pb}(\gamma,n)$ $-0.26$ $-0.10$
${}^{144}\mathrm{Sm}$ ${}^{142}\mathrm{Sm}(\gamma,n)$ $-0.15$ $-0.02$ ${}^{200}\mathrm{Pb}(\gamma,n)$ $0.17$ $0.07$
${}^{142}\mathrm{Sm}(\gamma,p)$ $-0.16$ $-0.02$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.24$ $0.10$
${}^{143}\mathrm{Sm}(\gamma,n)$ $-0.19$ $-0.03$
@@
${}^{146}\mathrm{Sm}(\gamma,\alpha)$ $-0.22$ $-0.03$
@@
${}^{150}\mathrm{Gd}(\gamma,\alpha)$ $-0.20$ $-0.03$
@@
Table C3: Correlations and $\zeta$ slopes between mass fraction and reaction rates for the 3×3$\times$3 ×3D-inspired mixing scenario.
@@
${}^{74}\mathrm{Se}$ ${}^{75}\mathrm{Se}(\gamma,n)$ $-0.81$ $-0.24$ ${}^{144}\mathrm{Sm}$ ${}^{142}\mathrm{Sm}(\gamma,n)$ $-0.23$ $-0.02$
${}^{78}\mathrm{Kr}$ ${}^{79}\mathrm{Kr}(\gamma,n)$ $-0.76$ $-0.20$ ${}^{142}\mathrm{Sm}(\gamma,p)$ $-0.23$ $-0.03$
${}^{84}\mathrm{Sr}$ ${}^{85}\mathrm{Sr}(\gamma,n)$ $-0.75$ $-0.26$ ${}^{143}\mathrm{Sm}(\gamma,n)$ $-0.30$ $-0.04$
${}^{92}\mathrm{Mo}$ ${}^{93}\mathrm{Mo}(\gamma,n)$ $-0.91$ $-0.23$ ${}^{146}\mathrm{Sm}(\gamma,\alpha)$ $-0.17$ $-0.02$
@@
${}^{96}\mathrm{Ru}$ ${}^{97}\mathrm{Ru}(\gamma,n)$ $-0.84$ $-0.25$ ${}^{150}\mathrm{Gd}(\gamma,\alpha)$ $-0.16$ $-0.02$
@@
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.23$ $-0.02$ ${}^{152}\mathrm{Gd}$ ${}^{151}\mathrm{Gd}(\gamma,n)$ $-0.17$ $-0.00$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $0.21$ $0.01$ ${}^{152}\mathrm{Dy}(\gamma,\alpha)$ $-0.39$ $-0.01$
${}^{102}\mathrm{Pd}$ ${}^{100}\mathrm{Pd}(\gamma,p)$ $-0.20$ $-0.06$ ${}^{154}\mathrm{Dy}(\gamma,\alpha)$ $-0.19$ $-0.00$
${}^{100}\mathrm{Pd}(\gamma,\alpha)$ $-0.22$ $-0.06$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $0.33$ $0.01$
${}^{103}\mathrm{Pd}(\gamma,n)$ $-0.71$ $-0.26$ ${}^{156}\mathrm{Dy}$ ${}^{157}\mathrm{Dy}(\gamma,n)$ $-0.21$ $-0.08$
${}^{106}\mathrm{Cd}$ ${}^{107}\mathrm{Cd}(\gamma,n)$ $-0.78$ $-0.27$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.18$ $-0.06$
@@
${}^{109}\mathrm{Cd}(\gamma,n)$ $-0.80$ $-0.24$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.15$ $0.05$
${}^{113}\mathrm{In}$ ${}^{110}\mathrm{Sn}(\gamma,p)$ $-0.22$ $-0.04$ ${}^{158}\mathrm{Dy}$ ${}^{157}\mathrm{Dy}(\gamma,n)$ $0.20$ $0.02$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.33$ $-0.06$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.19$ $-0.02$
@@
${}^{112}\mathrm{Sn}$ ${}^{113}\mathrm{Sn}(\gamma,n)$ $-0.78$ $-0.31$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.02$
${}^{114}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,p)$ $-0.17$ $-0.02$ ${}^{162}\mathrm{Er}$ ${}^{159}\mathrm{Er}(\gamma,n)$ $-0.29$ $-0.12$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.26$ $-0.03$ ${}^{159}\mathrm{Er}(\gamma,\alpha)$ $-0.17$ $-0.05$
${}^{113}\mathrm{Sn}(\gamma,n)$ $0.70$ $0.08$ ${}^{160}\mathrm{Er}(\gamma,n)$ $-0.22$ $-0.09$
${}^{122}\mathrm{Xe}(\gamma,p)$ $0.16$ $0.02$ ${}^{160}\mathrm{Er}(\gamma,\alpha)$ $-0.40$ $-0.13$
@@
${}^{115}\mathrm{Sn}$ ${}^{110}\mathrm{Sn}(\gamma,p)$ $-0.17$ $-0.02$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.27$ $0.10$
${}^{110}\mathrm{Sn}(\gamma,\alpha)$ $-0.25$ $-0.03$ ${}^{202}\mathrm{Pb}(\gamma,n)$ $0.15$ $0.06$
${}^{113}\mathrm{Sn}(\gamma,n)$ $0.73$ $0.08$ ${}^{164}\mathrm{Er}$ ${}^{164}\mathrm{Yb}(\gamma,n)$ $-0.24$ $-0.08$
${}^{122}\mathrm{Xe}(\gamma,\alpha)$ $0.24$ $0.03$ ${}^{164}\mathrm{Yb}(\gamma,\alpha)$ $-0.56$ $-0.24$
${}^{169}\mathrm{Lu}(\gamma,n)$ $0.15$ $0.01$ ${}^{165}\mathrm{Yb}(\gamma,n)$ $-0.17$ $-0.06$
${}^{120}\mathrm{Te}$ ${}^{121}\mathrm{Te}(\gamma,n)$ $-0.78$ $-0.24$ ${}^{196}\mathrm{Pb}(\gamma,n)$ $0.19$ $0.06$

### v2

@@ Appendix C Results of varying the input nuclear reactions
$^{74}\mathrm{Se}$ $^{75}\mathrm{Se}(\gamma,n)$ 0.83-0.83 0.22-0.22 $^{152}\mathrm{Gd}$ $^{150}\mathrm{Gd}(\gamma,\alpha)$ 0.18-0.18 0.00-0.00
$^{78}\mathrm{Kr}$ $^{79}\mathrm{Kr}(\gamma,n)$ 0.79-0.79 0.24-0.24 $^{151}\mathrm{Gd}(\gamma,n)$ 0.20-0.20 0.00-0.00
$^{84}\mathrm{Sr}$ $^{85}\mathrm{Sr}(\gamma,n)$ 0.80-0.80 0.34-0.34 $^{152}\mathrm{Dy}(\gamma,\alpha)$ 0.31-0.31 0.01-0.01
$^{92}\mathrm{Mo}$ $^{93}\mathrm{Mo}(\gamma,n)$ 0.91-0.91 0.14-0.14 $^{154}\mathrm{Dy}(\gamma,\alpha)$ 0.17-0.17 0.00-0.00
@@
$^{96}\mathrm{Ru}$ $^{97}\mathrm{Ru}(\gamma,n)$ 0.82-0.82 0.14-0.14 $^{196}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.00$
$^{100}\mathrm{Pd}(\gamma,\alpha)$ $0.21$ $0.04$ $^{156}\mathrm{Dy}$ $^{157}\mathrm{Dy}(\gamma,n)$ 0.28-0.28 0.12-0.12
$^{98}\mathrm{Ru}$ $^{97}\mathrm{Ru}(\gamma,n)$ $0.50$ $0.02$ $^{159}\mathrm{Er}(\gamma,n)$ 0.17-0.17 0.05-0.05
@@
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.31-0.31 0.01-0.01 $^{196}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.05$
@@
$^{102}\mathrm{Pd}$ $^{100}\mathrm{Pd}(\gamma,p)$ 0.18-0.18 0.04-0.04 $^{159}\mathrm{Dy}(\gamma,n)$ 0.21-0.21 0.03-0.03
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.21-0.21 0.04-0.04 $^{159}\mathrm{Er}(\gamma,n)$ 0.17-0.17 0.02-0.02
$^{103}\mathrm{Pd}(\gamma,n)$ 0.72-0.72 0.16-0.16 $^{160}\mathrm{Er}(\gamma,\alpha)$ $0.66$ $0.07$
$^{106}\mathrm{Cd}$ $^{107}\mathrm{Cd}(\gamma,n)$ 0.75-0.75 0.16-0.16 $^{196}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.02$
@@
$^{108}\mathrm{Cd}$ $^{107}\mathrm{Cd}(\gamma,n)$ $0.20$ $0.05$ $^{162}\mathrm{Er}$ $^{159}\mathrm{Er}(\gamma,n)$ 0.25-0.25 0.10-0.10
$^{109}\mathrm{Cd}(\gamma,n)$ 0.83-0.83 0.20-0.20 $^{159}\mathrm{Er}(\gamma,\alpha)$ 0.16-0.16 0.05-0.05
$^{113}\mathrm{In}$ $^{114}\mathrm{In}(\gamma,n)$ 0.32-0.32 0.05-0.05 $^{160}\mathrm{Er}(\gamma,n)$ 0.23-0.23 0.09-0.09
$^{110}\mathrm{Sn}(\gamma,p)$ 0.16-0.16 0.02-0.02 $^{160}\mathrm{Er}(\gamma,\alpha)$ 0.44-0.44 0.15-0.15
$^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.27-0.27 0.04-0.04 $^{161}\mathrm{Er}(\gamma,n)$ $0.16$ $0.06$
@@
$^{112}\mathrm{Sn}$ $^{113}\mathrm{Sn}(\gamma,n)$ 0.80-0.80 0.18-0.18 $^{202}\mathrm{Pb}(\gamma,n)$ $0.16$ $0.06$
$^{114}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.17-0.17 0.01-0.01 $^{164}\mathrm{Er}$ $^{164}\mathrm{Yb}(\gamma,n)$ 0.26-0.26 0.07-0.07
$^{113}\mathrm{Sn}(\gamma,n)$ $0.65$ $0.05$ $^{164}\mathrm{Yb}(\gamma,\alpha)$ 0.56-0.56 0.18-0.18
@@
$^{115}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.16-0.16 0.01-0.01 $^{168}\mathrm{Yb}$ $^{164}\mathrm{Yb}(\gamma,\alpha)$ 0.19-0.19 0.09-0.09
$^{113}\mathrm{Sn}(\gamma,n)$ $0.74$ $0.06$ $^{166}\mathrm{Yb}(\gamma,n)$ 0.22-0.22 0.08-0.08
$^{122}\mathrm{Xe}(\gamma,p)$ $0.16$ $0.01$ $^{166}\mathrm{Yb}(\gamma,\alpha)$ 0.38-0.38 0.13-0.13
@@
$^{120}\mathrm{Te}$ $^{121}\mathrm{Te}(\gamma,n)$ 0.76-0.76 0.19-0.19 $^{172}\mathrm{Hf}(\gamma,\alpha)$ $0.43$ $0.16$
@@
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.23-0.23 0.05-0.05 $^{202}\mathrm{Pb}(\gamma,n)$ $0.19$ $0.06$
$^{124}\mathrm{Xe}$ $^{122}\mathrm{Xe}(\gamma,p)$ 0.19-0.19 0.05-0.05 $^{174}\mathrm{Hf}$ $^{172}\mathrm{Hf}(\gamma,\alpha)$ 0.40-0.40 0.07-0.07
$^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.39-0.39 0.15-0.15 $^{174}\mathrm{W}(\gamma,\alpha)$ 0.19-0.19 0.04-0.04
$^{125}\mathrm{Xe}(\gamma,n)$ 0.54-0.54 0.23-0.23 $^{178}\mathrm{W}(\gamma,\alpha)$ $0.51$ $0.09$
$^{126}\mathrm{Xe}$ $^{122}\mathrm{Xe}(\gamma,\alpha)$ 0.35-0.35 0.14-0.14 $^{182}\mathrm{Os}(\gamma,\alpha)$ $0.18$ $0.03$
@@
$^{127}\mathrm{Xe}(\gamma,n)$ 0.63-0.63 0.27-0.27 $^{180}\mathrm{Ta}$ $^{179}\mathrm{Ta}(\gamma,n)$ 0.91-0.91 0.07-0.07
$^{130}\mathrm{Ba}$ $^{131}\mathrm{Ba}(\gamma,n)$ 0.82-0.82 0.13-0.13 $^{179}\mathrm{Ta}(\gamma,\alpha)$ 0.15-0.15 0.01-0.01
$^{132}\mathrm{Ba}$ $^{131}\mathrm{Ba}(\gamma,n)$ $0.36$ $0.14$ $^{180}\mathrm{W}$ $^{180}\mathrm{Os}(\gamma,n)$ 0.27-0.27 0.05-0.05
$^{133}\mathrm{Ba}(\gamma,n)$ 0.76-0.76 0.30-0.30 $^{180}\mathrm{Os}(\gamma,\alpha)$ 0.51-0.51 0.13-0.13
$^{138}\mathrm{La}$ $^{133}\mathrm{La}(\gamma,p)$ 0.20-0.20 0.07-0.07 $^{196}\mathrm{Pb}(\gamma,n)$ $0.23$ $0.04$
$^{135}\mathrm{La}(\gamma,n)$ 0.35-0.35 0.12-0.12 $^{184}\mathrm{Os}$ $^{182}\mathrm{Os}(\gamma,\alpha)$ 0.38-0.38 0.02-0.02
$^{136}\mathrm{La}(\gamma,n)$ $0.26$ $0.08$ $^{184}\mathrm{Pt}(\gamma,\alpha)$ 0.33-0.33 0.02-0.02
$^{137}\mathrm{La}(\gamma,n)$ 0.34-0.34 0.14-0.14 $^{186}\mathrm{Pt}(\gamma,\alpha)$ 0.15-0.15 0.01-0.01
$^{136}\mathrm{Ce}$ $^{137}\mathrm{Ce}(\gamma,n)$ 0.49-0.49 0.07-0.07 $^{188}\mathrm{Pt}(\gamma,\alpha)$ $0.23$ $0.01$
$^{138}\mathrm{Nd}(\gamma,n)$ 0.27-0.27 0.03-0.03 $^{196}\mathrm{Pb}(\gamma,n)$ $0.19$ $0.01$
$^{138}\mathrm{Nd}(\gamma,p)$ $0.48$ $0.06$ $^{190}\mathrm{Pt}$ $^{190}\mathrm{Hg}(\gamma,n)$ 0.30-0.30 0.02-0.02
$^{140}\mathrm{Nd}(\gamma,\alpha)$ $0.38$ $0.04$ $^{190}\mathrm{Hg}(\gamma,\alpha)$ 0.50-0.50 0.05-0.05
@@
$^{139}\mathrm{Ce}(\gamma,n)$ 0.52-0.52 0.05-0.05 $^{196}\mathrm{Hg}$ $^{196}\mathrm{Pb}(\gamma,n)$ 0.68-0.68 0.35-0.35
$^{138}\mathrm{Nd}(\gamma,p)$ $0.16$ $0.01$ $^{197}\mathrm{Pb}(\gamma,n)$ 0.26-0.26 0.10-0.10
$^{144}\mathrm{Sm}$ $^{142}\mathrm{Sm}(\gamma,n)$ 0.15-0.15 0.02-0.02 $^{200}\mathrm{Pb}(\gamma,n)$ $0.17$ $0.07$
$^{142}\mathrm{Sm}(\gamma,p)$ 0.16-0.16 0.02-0.02 $^{202}\mathrm{Pb}(\gamma,n)$ $0.24$ $0.10$
$^{143}\mathrm{Sm}(\gamma,n)$ 0.19-0.19 0.03-0.03
@@
$^{146}\mathrm{Sm}(\gamma,\alpha)$ 0.22-0.22 0.03-0.03
@@
$^{150}\mathrm{Gd}(\gamma,\alpha)$ 0.20-0.20 0.03-0.03
@@
Table C3: Correlations and $\zeta$ slopes between mass fraction and reaction rates for the $3\times$3D-inspired mixing scenario.
@@
$^{74}\mathrm{Se}$ $^{75}\mathrm{Se}(\gamma,n)$ 0.81-0.81 0.24-0.24 $^{144}\mathrm{Sm}$ $^{142}\mathrm{Sm}(\gamma,n)$ 0.23-0.23 0.02-0.02
$^{78}\mathrm{Kr}$ $^{79}\mathrm{Kr}(\gamma,n)$ 0.76-0.76 0.20-0.20 $^{142}\mathrm{Sm}(\gamma,p)$ 0.23-0.23 0.03-0.03
$^{84}\mathrm{Sr}$ $^{85}\mathrm{Sr}(\gamma,n)$ 0.75-0.75 0.26-0.26 $^{143}\mathrm{Sm}(\gamma,n)$ 0.30-0.30 0.04-0.04
$^{92}\mathrm{Mo}$ $^{93}\mathrm{Mo}(\gamma,n)$ 0.91-0.91 0.23-0.23 $^{146}\mathrm{Sm}(\gamma,\alpha)$ 0.17-0.17 0.02-0.02
@@
$^{96}\mathrm{Ru}$ $^{97}\mathrm{Ru}(\gamma,n)$ 0.84-0.84 0.25-0.25 $^{150}\mathrm{Gd}(\gamma,\alpha)$ 0.16-0.16 0.02-0.02
@@
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.23-0.23 0.02-0.02 $^{152}\mathrm{Gd}$ $^{151}\mathrm{Gd}(\gamma,n)$ 0.17-0.17 0.00-0.00
$^{110}\mathrm{Sn}(\gamma,\alpha)$ $0.21$ $0.01$ $^{152}\mathrm{Dy}(\gamma,\alpha)$ 0.39-0.39 0.01-0.01
$^{102}\mathrm{Pd}$ $^{100}\mathrm{Pd}(\gamma,p)$ 0.20-0.20 0.06-0.06 $^{154}\mathrm{Dy}(\gamma,\alpha)$ 0.19-0.19 0.00-0.00
$^{100}\mathrm{Pd}(\gamma,\alpha)$ 0.22-0.22 0.06-0.06 $^{160}\mathrm{Er}(\gamma,\alpha)$ $0.33$ $0.01$
$^{103}\mathrm{Pd}(\gamma,n)$ 0.71-0.71 0.26-0.26 $^{156}\mathrm{Dy}$ $^{157}\mathrm{Dy}(\gamma,n)$ 0.21-0.21 0.08-0.08
$^{106}\mathrm{Cd}$ $^{107}\mathrm{Cd}(\gamma,n)$ 0.78-0.78 0.27-0.27 $^{159}\mathrm{Er}(\gamma,n)$ 0.18-0.18 0.06-0.06
@@
$^{109}\mathrm{Cd}(\gamma,n)$ 0.80-0.80 0.24-0.24 $^{202}\mathrm{Pb}(\gamma,n)$ $0.15$ $0.05$
$^{113}\mathrm{In}$ $^{110}\mathrm{Sn}(\gamma,p)$ 0.22-0.22 0.04-0.04 $^{158}\mathrm{Dy}$ $^{157}\mathrm{Dy}(\gamma,n)$ $0.20$ $0.02$
$^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.33-0.33 0.06-0.06 $^{159}\mathrm{Er}(\gamma,n)$ 0.19-0.19 0.02-0.02
@@
$^{112}\mathrm{Sn}$ $^{113}\mathrm{Sn}(\gamma,n)$ 0.78-0.78 0.31-0.31 $^{202}\mathrm{Pb}(\gamma,n)$ $0.20$ $0.02$
$^{114}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,p)$ 0.17-0.17 0.02-0.02 $^{162}\mathrm{Er}$ $^{159}\mathrm{Er}(\gamma,n)$ 0.29-0.29 0.12-0.12
$^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.26-0.26 0.03-0.03 $^{159}\mathrm{Er}(\gamma,\alpha)$ 0.17-0.17 0.05-0.05
$^{113}\mathrm{Sn}(\gamma,n)$ $0.70$ $0.08$ $^{160}\mathrm{Er}(\gamma,n)$ 0.22-0.22 0.09-0.09
$^{122}\mathrm{Xe}(\gamma,p)$ $0.16$ $0.02$ $^{160}\mathrm{Er}(\gamma,\alpha)$ 0.40-0.40 0.13-0.13
@@
$^{115}\mathrm{Sn}$ $^{110}\mathrm{Sn}(\gamma,p)$ 0.17-0.17 0.02-0.02 $^{196}\mathrm{Pb}(\gamma,n)$ $0.27$ $0.10$
$^{110}\mathrm{Sn}(\gamma,\alpha)$ 0.25-0.25 0.03-0.03 $^{202}\mathrm{Pb}(\gamma,n)$ $0.15$ $0.06$
$^{113}\mathrm{Sn}(\gamma,n)$ $0.73$ $0.08$ $^{164}\mathrm{Er}$ $^{164}\mathrm{Yb}(\gamma,n)$ 0.24-0.24 0.08-0.08
$^{122}\mathrm{Xe}(\gamma,\alpha)$ $0.24$ $0.03$ $^{164}\mathrm{Yb}(\gamma,\alpha)$ 0.56-0.56 0.24-0.24
$^{169}\mathrm{Lu}(\gamma,n)$ $0.15$ $0.01$ $^{165}\mathrm{Yb}(\gamma,n)$ 0.17-0.17 0.06-0.06
$^{120}\mathrm{Te}$ $^{121}\mathrm{Te}(\gamma,n)$ 0.78-0.78 0.24-0.24 $^{196}\mathrm{Pb}(\gamma,n)$ $0.19$ $0.06$